# Understanding and Mitigating Memorization in Generative Models via Sharpness of Probability Landscapes

### v3

@@ D.2 Memorization Mitigation
  3. (Wen et al., 2024): As explained in Section D.1Wen et al. (2024) propose a differentiable metric based on the norm of the difference between the conditional and unconditional scores. Since memorized prompts empirically exhibit a large magnitude for this term, Wen et al. (2024) optimize the text embedding by directly minimizing it.
@@
Algorithm 2 provides a pseudo-code for SAIL algorithm. While Algorithm 2 shows the case of optimizing a single $\mathbf{x}_{T}$, in practice, it can simultaneously search for several memorization-free candidates by collectively optimizing several initializations in a batch fashion.
@@
Details of prompts in Figure 6.

### v4

@@ D.2 Memorization Mitigation
  3. (Wen et al., 2024): As explained in Section˜D.1Wen et al. (2024) propose a differentiable metric based on the norm of the difference between the conditional and unconditional scores. Since memorized prompts empirically exhibit a large magnitude for this term, Wen et al. (2024) optimize the text embedding by directly minimizing it.
@@
Algorithm˜2 provides a pseudo-code for SAIL algorithm. While Algorithm˜2 shows the case of optimizing a single $\mathbf{x}_{T}$, in practice, it can simultaneously search for several memorization-free candidates by collectively optimizing several initializations in a batch fashion.
@@
Details of prompts in Figure˜6.